OpenStudy (anonymous):

4/3(pi)r^3 =V I don't know how to make the symbols on here... solve for "r"

OpenStudy (anonymous):

Google that

OpenStudy (anonymous):

Its the volume of sphere

OpenStudy (anonymous):

It is the formula to figure the volume of a sphere. 4/3 x pi x r cubed pi = 3.1614 r = the radius of the sphere cubed means to multiply the radius by itself 3 times: r*r*r = r cubed. Exampled: What is the volume of a sphere with a radius of 3 inches? The radius = 3 inches, so r cubed = 3 x 3 x 3 = 9 so, 4/3 x 3.1416 x 9 = 37.7 cubic inches. Got it? Source: yahoo answer

OpenStudy (anonymous):

no, thanks though.. @Rohangrr

OpenStudy (anonymous):

http://www.mathway.com/problem.aspx?p=algebra

OpenStudy (anonymous):

\[\frac{ 4 }{ 3 }\pi r ^{3}=V \rightarrow r ^{3}=V \frac{ 3 }{ 4 \pi}\]

OpenStudy (anonymous):

then just get rid of the power by taking each side to the ^(1/3) power. Do you see what happened now?

OpenStudy (anonymous):

@ChmE thanks!